# The Relative Entropy of Expectation and Price

### v2

@@ Abstract
$p_{t}=-\frac{1}{\alpha}\log\mathbb{E}_{t}\exp[-\alpha p_{T}]$ (1)
for the price $p$ of a funded and hedged derivative security, where the price measure $\mathbb{E}$ has minimum entropy relative to economic expectations, and the parameter $\alpha$ is determined from the risk appetite of the investor. This price principle is easily applied to standard models for market evolution, with applications considered here including model risk analysis, deep hedging and decentralised finance.
@@ I Introduction
Suppose that two counterparties deploy initial capital $c_{1}>0$ and $c_{2}>0$ to the transaction and deposit the proceeds in liquid margin accounts with unit price $b_{1}>0$ and $b_{2}>0$ respectively. The first counterparty buys the derivative from the second counterparty, and the capital deployed must be sufficient to enter the transaction:
@@
With the proceeds and unused capital deposited in the respective margin accounts, the final margin available to settle the derivative is $(c_{1}-p)(1+db_{1}/b_{1})$ for the first counterparty and $(c_{2}+p)(1+db_{2}/b_{2})$ for the second counterparty. This is sufficient to settle the contractual commitments of the derivative when:
$-(c_{1}-p)(1+db_{1}/b_{1})\leq p+dp\leq(c_{2}+p)(1+db_{2}/b_{2})$ (3)
@@
$d\frac{c_{1}}{b_{1}}$ $=$ (7)
$\left(-\frac{c_{1}}{b_{1}}-d\frac{p}{b_{1}}\right)^{\!+}-\frac{b_{2}+db_{2}}{b_{1}+db_{1}}\left(d\frac{p}{b_{2}}-\frac{c_{2}}{b_{2}}\right)^{\!+}+d\frac{p}{b_{1}}$
$d\frac{c_{2}}{b_{2}}$ $=$
$\left(d\frac{p}{b_{2}}-\frac{c_{2}}{b_{2}}\right)^{\!+}-\frac{b_{1}+db_{1}}{b_{2}+db_{2}}\left(-\frac{c_{1}}{b_{1}}-d\frac{p}{b_{1}}\right)^{\!+}-d\frac{p}{b_{2}}$
demonstrating that the discounted return on capital matches the discounted return on the derivative adjusted by a put option on $d(p/b_{1})$ with strike $-c_{1}/b_{1}$, exercised when the first counterparty defaults, and a call option on $d(p/b_{2})$ with strike $c_{2}/b_{2}$, exercised when the second counterparty defaults. When assessing the viability of the trade, it is these net returns that must be considered, accounting for funding costs and the possibility that the settling counterparty defaults on its commitments.
This simple framework for bilateral margin is used in the following to derive equilibrium pricing expressions that account for funding and default, using a margin model based on the entropic risk metric – a convex risk metric that is chosen for its convenient decomposition in tiered derivative markets. It is an idealisation of real trading activity, and in particular does not account for the default protections provided by additional capital covering the net risks of the institution. Extensions to multilateral margin, which can be incorporated into the framework, are not considered here.
@@ II.1 Entropic risk metric
In this expression, the parameter $\alpha$ controls the risk adjustment, and is interpreted as the Lagrange multiplier for the variance constraint in a strategy that maximises the mean for a fixed variance. Incorporating higher moments, entropic-risk optimisation further extends the strategy by maximising the entropy-adjusted mean:
@@
Figure 4: Entropy-adjusted mean as a function of risk adjustment for three random variables. In each case, the variable is standardised by linear transformation to have zero mean and unit standard deviation. The entropy-adjusted mean is monotonic decreasing; it is unbounded for the normal variable, bounded below for the Poisson variable with ten expected jumps, and bounded above and below for the uniform variable.
@@ II.2 Optimising entropic risk
Risk adjustment is controlled by the parameter $\alpha$ in the range $0\leq\alpha\leq\infty$, and entropic risk is monotonic in this range:
@@ III Entropic pricing and hedging
The market of underlying securities accessible to the investor performs the two key functions of funding and hedging for the derivative security. Fix a finite investment horizon, and assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $q+dq$ and a derivative security with initial price $p$ and final price $p+dp$. The investor quantifies a priori the performance of the investment strategy using an initial assessment of the final economic state, encapsulated in the expectation measure $\mathbb{E}$ that maps variables at final time $t+dt$ to their expectations at initial time $t$. Following the strategy developed in the previous section, the optimal portfolio then maximises the entropy-adjusted mean return at a specified risk adjustment $\alpha$ related to the risk appetite of the investor.
@@
Underlying only: In this strategy, the investment comprises an underlying portfolio $\omega$ satisfying the initial condition: $0=\omega\cdot q$ (36) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[\omega\cdot dq]=$ (37) $-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha\omega\cdot dq]$ at the risk adjustment $\alpha>0$. The optimal portfolio $\omega$ that maximises this satisfies the calibration condition: $0=\mathbb{E}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]$ (38) where $r\,dt$ is the Lagrange multiplier used to implement the self-funding condition. Underlying and derivative: In this strategy, the investment comprises the derivative security and an underlying portfolio $(\omega-\delta)$ satisfying the initial condition: $0=p+(\omega-\delta)\cdot q$ (39) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[dp+(\omega-\delta)\cdot dq]=$ (40) $-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]$ at the risk adjustment $\alpha>0$. The hedge portfolio $\delta$ that maximises this satisfies the hedge condition: $0=\mathbb{E}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]]$ (41) where $(r+\alpha s)\,dt$ is the Lagrange multiplier used to implement the self-funding condition.
@@
Figure 7: Three examples of entropic hedging and pricing for a call option on an underlying security with return $dq$ modelled as a normal variable, a Poisson variable with ten expected jumps, and a uniform variable in the price measure. In each case, the return is standardised by linear transformation to have zero mean and unit standard deviation, and the derivative payoff for strike $k$ is $P:=(dq-k)^{+}$. Risk adjustment is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$. The graphs plot the delta and price per unit notional as a function of the strike for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Residual risk after hedging creates a bid-offer spread around the mid price and skews the optimal hedge ratio for positions with large notional.
@@
Figure 8: Three examples of entropic hedging and pricing for a digital option on an underlying security with return $dq$ modelled as a normal variable, a Poisson variable with ten expected jumps, and a uniform variable in the price measure. In each case, the return is standardised by linear transformation to have zero mean and unit standard deviation, and the derivative payoff for strike $k$ is $P:=1_{dq>k}$. Risk adjustment is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$. The graphs plot the delta and price per unit notional as a function of the strike for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Residual risk after hedging creates a bid-offer spread around the mid price and skews the optimal hedge ratio for positions with large notional.
@@
and the self-funding condition. Scaling the derivative notional is thus equivalent to scaling the risk appetite.
@@ III.5 Heston model
Figure 9: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 100 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimal entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a call option with unit payoff $P:=(Q-k)^{+}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk adjustment is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
@@
Figure 10: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 100 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimal entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a digital option with unit payoff $P:=1_{Q>k}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk adjustment is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
@@ V Entropic XVA
If the true costs of risk management are fully accounted in the return, reckless investments are naturally penalised by their associated margin costs. Revisiting the bilateral model outlined in the introduction, suppose that two counterparties deploy capital $c_{1}$ and $c_{2}$ to a transaction with initial price $p$, where $-c_{2}\leq p\leq c_{1}$. Excess capital is deposited in the margin account with price $b_{1}>0$ and $b_{2}>0$ respectively. The transaction is settled from the margin account with no additional funds injected, creating the risk that insufficient capital results in default by one of the counterparties.
Net of all considerations, the returns are:
$d\frac{c_{1}}{b_{1}}$ $=$ (84)
$\left(-\frac{c_{1}}{b_{1}}-d\frac{p}{b_{1}}\right)^{\!+}-\frac{b_{2}+db_{2}}{b_{1}+db_{1}}\left(d\frac{p}{b_{2}}-\frac{c_{2}}{b_{2}}\right)^{\!+}+d\frac{p}{b_{1}}$
$d\frac{c_{2}}{b_{2}}$ $=$
$\left(d\frac{p}{b_{2}}-\frac{c_{2}}{b_{2}}\right)^{\!+}-\frac{b_{1}+db_{1}}{b_{2}+db_{2}}\left(-\frac{c_{1}}{b_{1}}-d\frac{p}{b_{1}}\right)^{\!+}-d\frac{p}{b_{2}}$
To guarantee that default is avoided, the capital must exceed the maximum liability:
$c_{1}$ $\geq p-\operatorname*{ess\,inf}\!\left[\frac{P}{1+db_{1}/b_{1}}\right]$ (85)
$c_{2}$ $\geq\operatorname*{ess\,sup}\!\left[\frac{P}{1+db_{2}/b_{2}}\right]-p$
where $P:=p+dp$ is the final price. The first counterparty is at risk of default if the first condition is not met, and the second counterparty is at risk of default if the second condition is not met. These default risks need to be accounted for in the valuation.
@@
Viability of the transaction is assessed by the counterparty using their expectation measure $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$ and risk adjustment $\alpha_{1}$ and $\alpha_{2}$, with the threshold for execution given by the condition:
$\mathbb{E}_{1}[\alpha_{1}]\!\left[d\frac{c_{1}}{b_{1}}\right]$ $\geq 0$ (86)
$\mathbb{E}_{2}[\alpha_{2}]\!\left[d\frac{c_{2}}{b_{2}}\right]$ $\geq 0$
on the entropy-adjusted mean return on capital. To avoid the possibility of funding arbitrage, assume that the margin accounts have the same price, $b_{1}=b_{2}:=b$. With this assumption, the return simplifies:
$d\frac{c_{1}}{b}$ $=\frac{\bar{P}}{b+db}-\frac{p}{b}$ (87)
$d\frac{c_{2}}{b}$ $=\frac{p}{b}-\frac{\bar{P}}{b+db}$
where:
$\bar{P}$ ::=:= (88)
$\max[-(c_{1}-p)(1+db/b),\min[(c_{2}+p)(1+db/b),P]]$
is the final price floored such that the first counterparty defaults when the floor is breached and capped such that the second counterparty defaults when the cap is breached.
The returns for the counterparties are equal and opposite. This can only be viable if the counterparties have differing economic expectations and therefore attribute different values to the transaction. The return is monotonic in the initial price, and the threshold price $p_{1}$ and $p_{2}$ solves the implicit relationship:
$\frac{p_{1}}{b}$ $=\mathbb{E}_{1}[\alpha_{1}]\!\left[\frac{\bar{P}_{1}}{b+db}\right]$ (89)
$\frac{p_{2}}{b}$ $=\mathbb{E}_{2}[\alpha_{2}]\!\left[\frac{\bar{P}_{2}}{b+db}\right]$
@@
With these thresholds, the transaction is viable when the initial price $p$ is in the range $p_{2}\leq p\leq p_{1}$, and is therefore not viable when $p_{1}<p_{2}$. Price is thus constrained to the sector bounded by these viability thresholds, equilibrating to a price within the range. Since the entropy-adjusted mean is monotonic as a function of the notional, the range narrows as the notional increases and there is an equilibrium point at which no further transactions are viable. Equilibrium is thus discovered as the solution of the relationships:
$\frac{p}{b}$ $=\mathbb{E}_{1}[\lambda\alpha_{1}]\!\left[\frac{\bar{P}}{b+db}\right]$ (90)
$\frac{p}{b}$ $=\mathbb{E}_{2}[\lambda\alpha_{2}]\!\left[\frac{\bar{P}}{b+db}\right]$
@@
for the initial price per unit notional $p$ and the equilibrium notional $\lambda$. Terminating when the transaction exceeds the risk appetite of the investor, this procedure balances the differing expectations of the counterparties to identify their equilibrium price, accounting for their respective funding costs and risks of default.
@@
Figure 11: In this example, two counterparties have differing expectations for the future state. The first counterparty models the state as a uniform variable with zero mean and unit standard deviation. The second counterparty models the state as a normal variable with mean $\mu$ and unit standard deviation. The transaction is then a call option on the state with zero strike. In each graph, the entropy-adjusted mean return on capital, computed with risk adjustment $\alpha=1$, is plotted for both counterparties as a function of the option price per unit notional, including a range of values for the mean $\mu$ assumed by the second counterparty. The first counterparty is fully capitalised against default, and each row includes cases $c=\infty,0.5$ for the capital of the second counterparty. Each column includes cases $\lambda=0,1,2$ for the notional, where the case $\lambda=0$ corresponds to mid-pricing.
      The transaction is viable when the entropy-adjusted mean return on capital is positive for both counterparties. The first counterparty buys the option when its price is low, and the second counterparty sells the option when its price is high. The second counterparty is more likely to sell when the mean $\mu$ is lower, as this reduces the expected option payoff. Model risk increases as the notional $\lambda$ is increased, which pushes the threshold prices in opposite directions and makes the transaction less viable. Finally, for the second counterparty, decreasing the capital $c$ increases the risk of default. The CVA cost to the first counterparty and DVA benefit to the second counterparty then pushes both threshold prices lower.

### v3

@@ Abstract
$p=-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha P]$ (1)
for the price $p$ and payoff $P$ of a funded and hedged derivative security, where the price measure $\mathbb{E}$ has minimum entropy relative to economic expectations, and the parameter $\alpha$ matches the risk aversion of the investor. This price principle is easily applied to standard models for market evolution, with applications considered here including model risk analysis, deep hedging and decentralised finance.
@@ I Introduction
Suppose that two counterparties deploy initial capital $c_{1}>0$ and $c_{2}>0$ to the transaction and deposit the proceeds in a liquid margin account with positive unit price $b>0$. The first counterparty buys the derivative from the second counterparty, and the capital deployed must be sufficient to enter the transaction:
@@
With the proceeds and unused capital deposited in the margin account, the final margin available to settle the derivative is $(c_{1}-p)(1+db/b)$ for the first counterparty and $(c_{2}+p)(1+db/b)$ for the second counterparty. This is sufficient to settle the contractual commitments of the derivative when:
$-(c_{1}-p)(1+db/b)\leq p+dp\leq(c_{2}+p)(1+db/b)$ (3)
@@
$d\frac{c_{1}}{b}$ $=\left(-\frac{c_{1}}{b}-d\frac{p}{b}\right)^{\!+}-\left(d\frac{p}{b}-\frac{c_{2}}{b}\right)^{\!+}+d\frac{p}{b}$ (7)
$d\frac{c_{2}}{b}$ $=\left(d\frac{p}{b}-\frac{c_{2}}{b}\right)^{\!+}-\left(-\frac{c_{1}}{b}-d\frac{p}{b}\right)^{\!+}-d\frac{p}{b}$
demonstrating that the discounted return on capital matches the discounted return on the derivative adjusted by a put option on $d(p/b)$ with strike $-c_{1}/b$, exercised when the first counterparty defaults, and a call option on $d(p/b)$ with strike $c_{2}/b$, exercised when the second counterparty defaults. When assessing the viability of the trade, it is these net returns that must be considered, accounting for funding costs and the possibility that the settling counterparty defaults on its commitments.
This simple framework for bilateral margin is used in the following to derive equilibrium pricing expressions that account for funding and default. It is an idealisation of real trading activity, and in particular does not account for the default protections provided by additional capital covering the net risks of the institution. Extensions to multilateral margin, which can be incorporated into the framework, are not considered here.
The objective is to create foundations for pricing that do not rely on unrealistic assumptions on market completeness or the effectiveness of the trading strategy, properly account for residual model risk, and establish consistency across diverse applications including algorithmic trading, derivative pricing and hedging, and the modelling of margin and capital. In developing this objective, the entropic risk metric has dual application: as the determinant of margin used to protect against default; and as the a priori measure of performance used to optimise the investment strategy.
@@ II.1 Entropic risk metric
In this expression, the parameter $\alpha$ controls the risk aversion, and is interpreted as the Lagrange multiplier for the variance constraint in a strategy that maximises the mean for a fixed variance. Incorporating higher moments, entropic-risk optimisation further extends the strategy by maximising the entropy-adjusted mean:
@@
Figure 4: Entropy-adjusted mean as a function of risk aversion for three random variables. In each case, the variable is standardised by linear transformation to have zero mean and unit standard deviation. The entropy-adjusted mean is monotonic decreasing; it is unbounded for the normal variable, bounded below for the Poisson variable with ten expected jumps, and bounded above and below for the uniform variable.
@@ II.2 Optimising entropic risk
Risk aversion is controlled by the parameter $\alpha$ in the range $0\leq\alpha\leq\infty$, and entropic risk is monotonic in this range:
@@ III Entropic pricing and hedging
The market of underlying securities accessible to the investor performs the two key functions of funding and hedging for the derivative security. Fix a finite investment horizon, and assume the market is tiered into a set of underlying securities with initial price vector $q$ and final price vector $q+dq$ and a derivative security with initial price $p$ and final price $p+dp$. The investor quantifies a priori the performance of the investment strategy using an initial assessment of the final economic state, encapsulated in the expectation measure $\mathbb{E}$ that maps variables at final time $t+dt$ to their expectations at initial time $t$. Following the strategy developed in the previous section, the optimal portfolio then maximises the entropy-adjusted mean return at a specified risk aversion $\alpha$ for the investor.
@@
Underlying only: In this strategy, the investment comprises an underlying portfolio $\omega$ satisfying the initial condition: $0=\omega\cdot q$ (36) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[\omega\cdot dq]=$ (37) $-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha\omega\cdot dq]$ at the risk aversion $\alpha>0$. The optimal portfolio $\omega$ that maximises this satisfies the calibration condition: $0=\mathbb{E}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]$ (38) where $r\,dt$ is the Lagrange multiplier used to implement the self-funding condition. Underlying and derivative: In this strategy, the investment comprises the derivative security and an underlying portfolio $(\omega-\delta)$ satisfying the initial condition: $0=p+(\omega-\delta)\cdot q$ (39) so that the investment is self-funded. The return on the investment has entropy-adjusted mean: $\mathbb{E}[\alpha]$ $[dp+(\omega-\delta)\cdot dq]=$ (40) $-\frac{1}{\alpha}\log\mathbb{E}\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]$ at the risk aversion $\alpha>0$. The hedge portfolio $\delta$ that maximises this satisfies the hedge condition: $0=\mathbb{E}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(dp+(\omega-\delta)\cdot dq)]]$ (41) where $(r+\alpha s)\,dt$ is the Lagrange multiplier used to implement the self-funding condition.
@@
Figure 7: Three examples of entropic hedging and pricing for a call option on an underlying security with return $dq$ modelled as a normal variable, a Poisson variable with ten expected jumps, and a uniform variable in the price measure. In each case, the return is standardised by linear transformation to have zero mean and unit standard deviation, and the derivative payoff for strike $k$ is $P:=(dq-k)^{+}$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$. The graphs plot the delta and price per unit notional as a function of the strike for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Residual risk after hedging creates a bid-offer spread around the mid price and skews the optimal hedge ratio for positions with large notional.
@@
Figure 8: Three examples of entropic hedging and pricing for a digital option on an underlying security with return $dq$ modelled as a normal variable, a Poisson variable with ten expected jumps, and a uniform variable in the price measure. In each case, the return is standardised by linear transformation to have zero mean and unit standard deviation, and the derivative payoff for strike $k$ is $P:=1_{dq>k}$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$. The graphs plot the delta and price per unit notional as a function of the strike for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Residual risk after hedging creates a bid-offer spread around the mid price and skews the optimal hedge ratio for positions with large notional.
@@
and the self-funding condition. Scaling the derivative notional is thus equivalent to scaling the risk aversion.
@@ III.5 Heston model
Figure 9: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 100 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a call option with unit payoff $P:=(Q-k)^{+}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
@@
Figure 10: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 100 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a digital option with unit payoff $P:=1_{Q>k}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 11: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 10,000 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a call option with unit payoff $P:=(Q-k)^{+}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 12: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 10,000 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a digital option with unit payoff $P:=1_{Q>k}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
@@ V Entropic XVA
Figure 13: In this example, the trader receives the initial derivative price $p$ and deploys additional capital $c$ to guarantee settlement, with $c+p\geq 0$. These proceeds are deposited in a margin account with unit price $b>0$. The final derivative price $p+dp$ is settled up to the value $(c+p)(1+db/b)$ of the margin account, with any remainder returned to the trader. Default occurs when the margin account is insufficient to meet the contractual obligation. In the expression for the final settlement, the cap is denoted by the superscript.
Figure 14: Two counterparties deploy capital $c_{1}$ and $c_{2}$ respectively as their guarantee of settlement, with $c_{1}-p\geq 0$ and $c_{2}+p\geq 0$, and they default on their contractual obligation if this is insufficient. Net of all considerations, the counterparties share the combined proceeds from the margin accounts, and the final return on capital is floored and capped by the respective default events. In the expressions for the final settlement and capital, the floor is denoted by the subscript and the cap is denoted by the superscript.
Figure 15: Eliminating internal settlements, the trade is simplified as a funding pool with contributions from both counterparties shared according to the contractual settlement of the derivative security adjusted for default. Using the entropic margin model, these settlements can be encoded in a smart contract for decentralised distribution.
@@
If the true costs of risk management are fully accounted in the return, reckless investments are naturally penalised by their associated margin costs. Revisiting the bilateral model outlined in the introduction, suppose that two counterparties deploy capital $c_{1}$ and $c_{2}$ to a trade with initial price $p$. Excess capital is deposited in the margin account with unit price $b>0$. The trade is settled from the margin account with no additional funds injected, creating the risk that insufficient capital results in default by one of the counterparties.
As shown in Figures 15, 15 and 15, net of all considerations the capital returned from the trade is:
$c_{1}+dc_{1}$ $=(b+db)\!\left(\frac{c_{1}}{b}+d\frac{p}{b}\right)_{0}^{(c_{1}+c_{2})/b}$ (84)
$=c_{1}(1+db/b)+\left(dp-\frac{p}{b}db\right)_{-c_{1}(1+db/b)}^{c_{2}(1+db/b)}$
$c_{2}+dc_{2}$ $=(b+db)\!\left(\frac{c_{2}}{b}-d\frac{p}{b}\right)_{0}^{(c_{1}+c_{2})/b}$
$=c_{2}(1+db/b)-\left(dp-\frac{p}{b}db\right)_{-c_{1}(1+db/b)}^{c_{2}(1+db/b)}$
where the subscript denotes the floor and the superscript denotes the cap. The combined proceeds from the margin accounts are shared between the counterparties, adjusting for the funded return on the derivative security, floored and capped so that the capital is positive.
The first counterparty defaults when the floor is breached and the second counterparty defaults when the cap is breached. Each counterparty assesses the threshold for default using their expectation measure $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$, and default is avoided when the capital exceeds the maximum liability:
$c_{1}$ $\geq p-\operatorname*{ess\,inf}\nolimits_{1}\!\left[\frac{P}{1+db/b}\right]$ (85)
$c_{2}$ $\geq\operatorname*{ess\,sup}\nolimits_{2}\!\left[\frac{P}{1+db/b}\right]-p$
where $P:=p+dp$ is the final price. Setting these inequalities to equalities generates the minimum capital required to fully guarantee settlement, under the assumptions of the respective expectation measures. More generally, the counterparties can exchange certainty of settlement for a reduction in capital. In the entropic margin model, this is parametrised by the default aversion $\beta_{1}$ and $\beta_{2}$ with:
$m_{1}$ $:=c_{1}-p=-\mathbb{E}_{1}[\beta_{1}]\!\left[\frac{P}{1+db/b}\right]$ (86)
$m_{2}$ $:=c_{2}+p=\mathbb{E}_{2}[-\beta_{2}]\!\left[\frac{P}{1+db/b}\right]$
where the case $\beta_{1}=\infty$ means the first counterparty provides full default protection and the case $\beta_{2}=\infty$ means the second counterparty provides full default protection.
@@
Viability of the trade is assessed by the counterparty using their risk aversion $\alpha_{1}$ and $\alpha_{2}$. Consider the threshold for execution given by the condition:
$\mathbb{E}_{1}[\alpha_{1}]\!\left[\frac{c_{1}+dc_{1}}{1+db/b}\right]$ $\geq c_{1}$ (87)
$\mathbb{E}_{2}[\alpha_{2}]\!\left[\frac{c_{2}+dc_{2}}{1+db/b}\right]$ $\geq c_{2}$
on the entropy-adjusted mean return on funded capital. These conditions impose bounds on the viable price:
$p$ $\leq p_{1}:=\mathbb{E}_{1}[\alpha_{1}]\!\left[\left(\frac{P}{1+db/b}\right)_{\!-m_{1}}^{\!m_{2}}\right]$ (88)
$p$ $\geq p_{2}:=\mathbb{E}_{2}[-\alpha_{2}]\!\left[\left(\frac{P}{1+db/b}\right)_{\!-m_{1}}^{\!m_{2}}\right]$
@@
The trade is viable when the initial price $p$ is in the range $p_{2}\leq p\leq p_{1}$, and is therefore not viable when $p_{1}<p_{2}$. Price is thus constrained to the sector bounded by these viability thresholds, equilibrating to a price within the range. Since the entropy-adjusted mean is monotonic as a function of the notional, the range narrows as the notional increases and there is an equilibrium point at which no further trades are viable. Equilibrium is thus discovered as the solution of the relationship:
[…]
@@
for the initial price per unit notional $p$ and the equilibrium notional $\lambda$. Terminating when the trade exceeds the risk appetite of the investor, this procedure balances the differing expectations of the counterparties to identify their equilibrium price, accounting for their respective funding costs and risks of default.
@@
Figure 16: In this example, two counterparties have differing expectations for the future state. The first counterparty models the state as a uniform variable with zero mean and unit standard deviation. The second counterparty models the state as a normal variable with mean $\mu$ and unit standard deviation. The trade is then a call option on the state with zero strike. In each graph, the entropy-adjusted mean return on capital, computed with risk aversion $\alpha=1$, is plotted for both counterparties as a function of the option price per unit notional, including a range of values for the mean $\mu$ assumed by the second counterparty. The first counterparty is fully capitalised against default, and each row includes cases $c=\infty,0.5$ for the capital of the second counterparty. Each column includes cases $\lambda=0,1,2$ for the notional, where the case $\lambda=0$ corresponds to mid-pricing.
      The trade is viable when the entropy-adjusted mean return on capital is positive for both counterparties. The first counterparty buys the option when its price is low, and the second counterparty sells the option when its price is high. The second counterparty is more likely to sell when the mean $\mu$ is lower, as this reduces the expected option payoff. Model risk increases as the notional $\lambda$ is increased, which pushes the threshold prices in opposite directions and makes the trade less viable. Finally, for the second counterparty, decreasing the capital $c$ increases the risk of default. The CVA cost to the first counterparty and DVA benefit to the second counterparty then pushes both threshold prices lower.